# Finite-Time Behavior of Erlang-C Model: Mixing Time, Mean Queue Length and Tail Bounds

### v3

@@ Abstract.
Service systems like data centers and ride-hailing are popularly modeled as queueing systems in the literature. Such systems are primarily studied in the steady state due to their analytical tractability. However, almost all applications in real life do not operate in a steady state, so there is a clear discrepancy in translating theoretical queueing results to practical applications. To this end, we provide a finite-time convergence for Erlang-C systems (also known as $M/M/n$ queues), providing a stepping stone towards understanding the transient behavior of more general queueing systems. We obtain a bound on the Chi-square distance between the finite time queue length distribution and the stationary distribution for a finite number of servers. We then use these bounds to study the behavior in the many-server heavy-traffic asymptotic regimes. The Erlang-C model exhibits a phase transition at the so-called Halfin-Whitt regime. We show that our mixing rate matches the limiting behavior in the Super-Halfin-Whitt regime and matches up to a constant factor in the Sub-Halfin-Whitt regime.
@@
To prove such a result, we employ the Lyapunov-Poincaré approach, where we first carefully design a Lyapunov function to obtain a negative drift outside a finite set. Within the finite set, we develop different strategies depending on the properties of the finite set to get a handle on the mixing behavior via a local Poincaré inequality. A key aspect of our methodological contribution is in obtaining tight guarantees in these two regions, which when combined, give us tight mixing time bounds. We believe that this approach is of independent interest for studying mixing in reversible countable-state Markov chains more generally.
Queuing Theory, Drift Method, Markov Chain mixing
copyright: acmlicensedjournalyear: 2025copyright: rightsretainedconference: Abstracts of the 2025 ACM SIGMETRICS International Conference on Measurement and Modeling of Computer Systems; June 9–13, 2025; Stony Brook, NY, USA.booktitle: Abstracts of the 2025 ACM SIGMETRICS International Conference on Measurement and Modeling of Computer Systems (SIGMETRICS Abstracts ’25), June 9–13, 2025, Stony Brook, NY, USAisbn: 979-8-4007-1593-8/25/06doi: 10.1145/3726854.3727287journal: JACMjournalvolume: 37journalnumber: 4article: 111publicationmonth: 8ccs: Mathematics of computing Probability and statisticsccs: Applied computing Operations research
1. Introduction
@@
Queueing theory has proved to be a successful tool for studying several service systems like analyzing data centers (Maguluri and Srikant, 2014; Maguluri et al., 2014, 2012), wireless networks (Nguyen and Modiano, 2023, 2024a, 2024b), and ride-hailing and EV systems (Varma et al., 2021, 2023b, 2023a). A popular approach is to study such systems in the steady-state owing to the tractability of analysis. However, these systems rarely operate in a steady state and so, the theoretical analysis does not directly translate into real-life performance guarantees. Furthermore, a typical goal in the performance analysis of service systems is to understand queue length/waiting time behavior. However, it is not possible to obtain the exact queue length distribution except in special cases such as the $M/M/1$ queue (Robert, 2003; Leguesdron et al., 1993; Morse, 1955). So typically, one uses the steady-state behavior as a proxy for the finite-time behavior. However, this approximation is useful only when the system mixes quickly, i.e. approaches the steady-state behavior quickly. These discrepancies motivate us to analyze the rate of convergence to the steady state. Such a result allows one to better understand the applicability of the steady-state results. As the transient analysis of queueing systems is known to be challenging, we restrict our attention to one of the simplest systems, the Erlang-C model (also known as the $M/M/n$ queue).
@@
Moreover, sometimes it is not even possible to obtain the closed form steady-state distribution of the system in many queuing systems, and therefore one resorts to asymptotic analysis such as heavy traffic analysis (Hurtado-Lange and Maguluri, 2020). Recent works have characterized the rate of convergence in these regimes, thus obtaining a handle on the pre-asymptotic queue length behavior (Jhunjhunwala et al., 2023; Braverman et al., 2017). Since these approximation errors are larger when one is further away from the asymptotic regime, and therefore one has a better handle on the stationary distribution when one is closer to the asymptotic regime.
@@
However, prior works suggest that the queuing systems mix slower as one gets closer to the asymptotic regimes, for example, in the Heavy Traffic regime (Pang et al., 2007; Robert, 2003; Gamarnik and Goldberg, 2013). And so, we have a dilemma where the system has a larger steady-state approximation error if it is further away from the asymptotic regime, but mixes slower when it is closer to the asymptotic regime. And so, it is important to characterize the goodness of the steady-state analysis and asymptotic analysis to the finite-time queue length behavior. Therefore, the goal of this paper is to develop a theoretical framework to analyze the mixing time of reversible Markov chains, which entails Erlang-C systems. Having such a framework would allow us to obtain queue length bounds in finite-time and pre-asymptotic regimes (i.e. with a finite number of servers) for the $M/M/n$ queue and finite-time analysis for other various applications such as learning policies for queues (van Kempen et al., 2024).
1.1. Contributions
@@
where the $\square$ term is called the mixing rate of the system and $\chi$ is the square root of the Chi-square distance, which would be defined in Subsection 2.2. Our results are true for all time $t$ and for all $n$ (number of servers). Furthermore, our results are tight in the sense that as $n$ goes to infinity, our mixing rate matches the mixing rate of $M/M/1$ and $M/M/\infty$, depending on the regime. In particular, for $\alpha\in(1/2,\infty)$, the system behavior resembles an $M/M/1$ system with service rate $n$ as most of the time all $n$ servers will be busy (i.e. the queue length is $n$ or above). On the other hand, for $\alpha\in(0,1/2)$, the system will have an idle server and resembles $M/M/\infty$. Therefore, the mixing time behavior of these systems will have
We obtain finite-time convergence results for the following regimes:
  • Super-Halfin-Whitt regime ($\alpha\in(1/2,\infty)$): In this regime, we show the convergence rate to stationarity to be $e^{-C_{n}(\sqrt{n}-\sqrt{\lambda_{n}})^{2}t}$ where $C_{n}$ is a positive parameter such that $\lim_{n\rightarrow\infty}C_{n}=1$ for $\alpha\in(1/2,1)$ and $C_{n}=1$ for $\alpha\in[1,\infty)$. In particular, our result for $\alpha\in[1,\infty)$ recovers the bound in (Zeifman, 1991). Moreover, to the best of our knowledge, our result in the regime $\alpha\in(1/2,1)$ is new in the literature and is a contribution of this paper, covering the gap $\alpha\in(1/2,1)$ between (Gamarnik and Goldberg, 2013; Zeifman, 1991). Our results are generalizations of the finite-time behavior of $M/M/1$ since they match the limiting behavior of $M/M/n$ in this regime, which is expected to resemble the behavior of $M/M/1$ with a service rate $n$ (whose mixing rate is known to be $(\sqrt{n}-\sqrt{\lambda_{n}})^{2}$ (Morse, 1955; Van Doorn, 1985; Robert, 2003; Gamarnik and Goldberg, 2013)).
  • Sub-Halfin-Whitt regime ($\alpha\in(0,1/2)$): We show that the mixing rate of $M/M/n$ queue in the Sub-Halfin-Whitt regime is $D_{n}$, where $D_{n}$ is a positive parameter satisfying $\lim_{n\rightarrow\infty}D_{n}=1/25$, which matches the asymptotic behavior up to a constant factor. To the best of our knowledge, we are the first to achieve a constant mixing time guarantee for $M/M/n$ in the Sub-Halfin-Whitt regime. Moreover, if we have $\lambda_{n}\in\mathbb{Z}^{+}$ then we show that one can match the asymptotic behavior at the limit, i.e. the mixing rate approaches $1$ as in the $M/M/\infty$ system (Chafaï, 2006). We conjecture that the condition $\lambda_{n}\in\mathbb{Z}$ is not required, and getting a mixing rate approaching $1$ without this condition will be a future work.
  • Mean Field regime $\left(\lim_{n\rightarrow\infty}\frac{\lambda_{n}}{n}=c\in[0,1)\right)$: Finally, in this regime, we show the mixing rate of the system approaches $1-\sqrt{c}$ as $n$ goes to infinity. When $c=0$, we have the mixing approaches $1$, which matches the mixing rate of $M/M/\infty$ at the asymptotic. Notably, this result does not rely on the arrival rate being an integer to match the mixing rate of $M/M/\infty$ at the asymptotic.
@@
  • Finite-time concentration bound of the number of customers: we obtain a finite-time concentration bound for the queue length. Our tail bound results match the steady-state and asymptotic behavior in (Jhunjhunwala et al., 2023; Braverman et al., 2017). The queue length concentration bound results are formally stated in Corollary 5.
@@
  • Probability of having an idle server: We obtain a bound on the probability of having an idle server at time $t$ (i.e. the queue length is less than $n$). Moreover, we also recover that as $n$ goes to infinity, this probability goes to 00 in the Super-Halfin-Whitt regime and goes to a value close to $1$ in the Sub-Halfin-Whitt regime. The full result is stated in Corollary 6.
@@
In summary, our finite-time behavior characterizations match the steady-state and asymptotic results previously established in (Jhunjhunwala et al., 2023; Braverman et al., 2017) and provide a holistic insight for Erlang-C systems.
Regime Mean Field Sub-HW Halfin-Whitt Super-HW
@@
$\lambda_{n}/n\rightarrow c$ $\lambda_{n}=n-B\sqrt{n}$
$\mathbb{E}\left[[q-n]^{+}\right]$ $O(1)$ $O(\sqrt{n})$ $O(n^{\alpha})$
Tail behav. Gaussian Exponential
Mixing rate $L_{n}$ $D_{n}$ $1-o(1)$ $\frac{B^{2}}{4}$ $C_{n}(\sqrt{n}-\sqrt{\lambda_{n}})^{2}$ $(\sqrt{n}-\sqrt{\lambda_{n}})^{2}$
$e^{-\square t}$ $L_{n}\rightarrow 1-\sqrt{c}$ $D_{n}\rightarrow\frac{1}{25}$ $C_{n}\rightarrow 1$
Ours (Gamarnik and Goldberg, 2013) Ours Ours, (Zeifman, 1991)
@@
Table 1. Overview of mixing time analysis for $M/M/n$ for $\lambda_{n}=n-n^{1-\alpha},\mu=1$ where $\alpha$ is a parameter. The number of waiting jobs is denoted as $\left[q-n\right]^{+}=\max\{q-n,0\}$, and the asymptotic behavior is the queue length distribution behavior as $n\rightarrow\infty$. When $\alpha\approx\infty$, this is called the Classical Heavy Traffic regime. For the Mean Field regime, we write $\alpha\approx 0$ as an intuitive explanation, whereas the actual condition is $\lim_{n\rightarrow\infty}\frac{\lambda_{n}}{n}=c\in[0,1)$.
1.2. Our approach: The Lyapunov-Poincaré method
It is well-known that one can obtain exponential ergodicity for a stochastic system if the system admits some Lyapunov drift outside of some finite set (Meyn and Tweedie, 1994; Down et al., 1995; Rosenthal, 1995; Lund et al., 1996; Baxendale, 2005; Douc et al., 2018; Meyn and Tweedie, 2012; Raginsky et al., 2017; Taghvaei and Mehta, 2022; Andrieu et al., 2022). In contrast, our goal is to get a handle on the mixing rate. While one can easily get a handle on the mixing outside of the said set using the negative drift, it is much harder to handle the behavior inside the finite set. The special case when the finite set is a singleton (which is the case in an $M/M/1$ queue) is well understood (Lund et al., 1996). When the finite set is not a singleton, one typically uses a ”stitching theorem” to combine knowledge of the drift outside the finite set and local mixing behavior inside the finite set to obtain a mixing time bound for the entire system. Prior work has explored the use of minorization (Down et al., 1995; Rosenthal, 1995; Baxendale, 2005; Taghvaei and Mehta, 2022), contraction (Durmus and Moulines, 2014; Hairer et al., 2011) within the finite set or obtain the local Poincaré constant by using the radius of the finite set (Dominique et al., 2008; Raginsky et al., 2017; Divol et al., 2024), and then a corresponding stitching theorem to obtain mixing rates. However, these approaches either do not lead to a tight mixing rate inside the finite set (especially for high-dimensional settings (Qin and Hobert, 2020)) or are hard to obtain in practice. To this end, we propose the Lyapunov-Poincaré method, where we combine the Lyapunov drift outside of a finite set with a local Poincaré inequality inside the finite set to establish mixing results. Our work extends the previously established Lyapunov drift framework (Meyn and Tweedie, 1994; Down et al., 1995; Rosenthal, 1995; Lund et al., 1996; Baxendale, 2005; Douc et al., 2018; Meyn and Tweedie, 2012; Raginsky et al., 2017; Taghvaei and Mehta, 2022; Andrieu et al., 2022) to allow a more fine-grained mixing analysis, which eventually enables us to obtain tight mixing bounds. In particular, we develop new stitching theorems so that depending on the properties of the finite set, we can better combine the drift properties outside the finite set with the local mixing behavior inside the finite set (which is characterized in terms of local Poincaré inequalities) so as to obtain more fine-grained mixing analysis.
  • Local canonical path method: The canonical path method, previously developed in (Jerrum and Sinclair, 1989), is a powerful technique used to analyze the mixing time of finite discrete Markov chains by bounding how ”congested” the state space is with respect to moving from one state to another. In particular, the canonical path method estimates the mixing time of a finite discrete Markov chain with the congestion ratio, which scales with the stationary distribution and the sum of the path length of paths passing through a particular edge. As such, the canonical path method can be viewed as a conductance-typed approach to obtain mixing. Notably, the canonical path method obtains a tight mixing bound when the Markov chain is a symmetric random walk on a path graph (Berestycki, 2020; Levin and Peres, 2017), which is also the behavior of the Erlang-C system in the finite set when an appropriate Lyapunov function is chosen. By applying the local canonical path method, we obtain a tight weighted local Poincare inequality of the Erlang-C system inside the finite set, which allows us to use our stitching theorems to obtain the mixing rate bound for the entire system. We formally state our local canonical path lemma in Lemma 1.
  • Truncation: On the other hand, the aforementioned local canonical path method is only viable when the distribution inside the finite set is roughly uniform. For other finite sets whose distributions are not necessarily uniform, we propose the truncation method to obtain local mixing results inside such sets. In essence, if we can find another system with known Poincaré constant, which behaves identically inside the finite set, then we can use it to get a handle on the local Poincaré constant of the original system. With this method, we obtain tight mixing time bounds up to a constant factor for $M/M/n$ queues in the Mean Field regime, i.e. when $\lim_{n\rightarrow\infty}\frac{\lambda_{n}}{n}=c\in(0,1)$. For $c=0$ (the so-called Light Traffic regime), we have the mixing rate approaches to $1$, which matches the limiting behavior of the system.
1.3. Related works

### v4

# Finite-Time behaviors of M/M/$n$ Queue: Mixing Time, Mean Queue Length and Tail Bounds

@@ Abstract
Service systems like data centers and ride-hailing are popularly modeled as queueing systems in the literature. Such systems are primarily studied in the steady state due to their analytical tractability. However, almost all applications in real life do not operate in a steady state, so there is a clear discrepancy in translating theoretical queueing results to practical applications. To this end, we provide a finite-time convergence for Erlang-C systems (also known as $M/M/n$ queues), providing a stepping stone towards understanding the transient behavior of more general queueing systems. We obtain a bound on the Chi-square distance between the finite time queue length distribution and the stationary distribution for a finite number of servers. We then use these bounds to study the behavior in the many-server heavy-traffic asymptotic regimes. The Erlang-C model exhibits a phase transition at the so-called Halfin-Whitt regime. We show that our mixing rate matches the limiting behavior in the Super-Halfin-Whitt regime, and matches up to a constant factor in the Sub-Halfin-Whitt regime.
@@
To prove such a result, we employ the Lyapunov-Poincaré approach, where we first carefully design a Lyapunov function to obtain a negative drift outside a finite set. Within the finite set, we develop different strategies depending on the properties of the finite set to get a handle on the mixing behavior via a local Poincaré inequality. A key aspect of our methodological contribution is in obtaining tight guarantees in these two regions, which when combined give us tight mixing time bounds. We believe that this approach is of independent interest for studying mixing in reversible countable-state Markov chains more generally.
1 Introduction
@@
Queueing theory has proved to be a successful tool for studying several service systems like analyzing data centers [46, 48, 47], wireless networks [58, 59, 60], and ride-hailing and EV systems [81, 83, 82]. A popular approach is to study such systems in the steady-state owing to the tractability of analysis. However, these systems rarely operate in a steady state and so, the theoretical analysis does not directly translate into real-life performance guarantees. Furthermore, a typical goal in the performance analysis of service systems is to understand queue length/waiting time behavior. However, it is not possible to obtain the exact queue length distribution except in special cases such as the $M/M/1$ queue [67, 39, 54]. So typically, one uses the steady-state behavior as a proxy for the finite-time behavior. However, this approximation is useful only when the system mixes quickly, i.e. approaches the steady-state behavior quickly. These discrepancies motivate us to analyze the rate of convergence to the steady state. Such a result allows one to better understand the applicability of the steady-state results. As the transient analysis of queueing systems is known to be challenging, we restrict our attention to one of the simplest systems, the Erlang-C model (also known as the $M/M/n$ queue).
@@
Moreover, sometimes it is not even possible to obtain the closed form steady-state distribution of the system in many queuing systems, and therefore one resorts to asymptotic analysis such as heavy traffic analysis [33]. Recent works have characterized the rate of convergence in these regimes, thus obtaining a handle on the pre-asymptotic queue length behavior [35, 9]. Since these approximation errors are larger when one is further away from the asymptotic regime, and therefore one has a better handle on the stationary distribution when one is closer to the asymptotic regime.
@@
However, prior works suggest that the queuing systems mix slower as one gets closer to the asymptotic regimes, for example, in the Heavy Traffic regime [61, 67, 29]. And so, we have a dilemma where the system has a larger steady-state approximation error if it is further away from the asymptotic regime, but mixes slower when it is closer to the asymptotic regime. And so, it is important to characterize the goodness of the steady-state analysis and asymptotic analysis to the finite-time queue length behavior. Therefore, the goal of this paper is to develop a theoretical framework to analyze the mixing time of reversible Markov chains, which entails Erlang-C systems. Having such a framework would allow us to obtain queue length bounds in finite-time and pre-asymptotic regimes (i.e. with a finite number of servers) for the $M/M/n$ queue and finite-time analysis for other various applications such as learning policies for queues [79] or Markov chain perturbations [52, 69].
1.1 Contributions
@@
where the $\square$ term is called the mixing rate of the system and $\chi$ is the square root of the Chi-square distance, which would be defined in Subsection 2.2. Our results are true for all time $t$ and for all $n$ (number of servers) and such finite-time results can be obtained from any initial queue length distribution. Furthermore, our results are tight in the sense that as $n$ goes to infinity, our mixing rate matches the mixing rate of $M/M/1$ and $M/M/\infty$, depending on the regime. In particular, for $\alpha\in(1/2,\infty)$, the system behavior resembles an $M/M/1$ system with service rate $n$. On the other hand, for $\alpha\in(0,1/2)$, the system will have an idle server with a high probability, and this resembles the behavior of an $M/M/\infty$ system.
We describe our finite-time convergence results for each regime in detail as follows.
  • Super-Halfin-Whitt regime ($\alpha\in(1/2,\infty)$): In this regime, we show the convergence rate to stationarity to be $e^{-C_{n}(\sqrt{n}-\sqrt{\lambda_{n}})^{2}t}$ where $C_{n}$ is a positive parameter such that $\lim_{n\rightarrow\infty}C_{n}=1$ for $\alpha\in(1/2,1)$ and $C_{n}=1$ for $\alpha\in[1,\infty)$. Our results are generalizations of the finite-time behavior of $M/M/1$ since they match the limiting behavior of $M/M/n$ in this regime, which is expected to resemble the behavior of $M/M/1$ with a service rate $n$ (whose mixing rate is known to be $(\sqrt{n}-\sqrt{\lambda_{n}})^{2}$ [54, 73, 67, 29]).
  • Sub-Halfin-Whitt regime ($\alpha\in(0,1/2)$): We show that the mixing rate of $M/M/n$ queue in the Sub-Halfin-Whitt regime is $D_{n}$, where $D_{n}$ is a positive parameter satisfying $\lim_{n\rightarrow\infty}D_{n}=1/25$, which matches the asymptotic behavior up to a constant factor. Moreover, if we have $\lambda_{n}\in\mathbb{Z}^{+}$ then we show that one can match the asymptotic behavior at the limit, i.e. the mixing rate approaches $1$ as in the $M/M/\infty$ system [12]. We conjecture that the condition $\lambda_{n}\in\mathbb{Z}$ is not required, and getting a mixing rate approaching $1$ without this condition will be a future work.
  • Mean Field regime $\left(\lim_{n\rightarrow\infty}\frac{\lambda_{n}}{n}=c\in[0,1)\right)$: Finally, in this regime, we show the mixing rate of the system approaches 1c1-c as $n$ goes to infinity. When $c=0$, we have the mixing rate approaches $1$, which matches the mixing rate of $M/M/\infty$ at the asymptotic. Notably, this result does not rely on the arrival rate being an integer to match the mixing rate of $M/M/\infty$ at the asymptotic.
@@
  • Finite-time concentration bound of the number of customers: we obtain a finite-time concentration bound for the queue length. Our tail bound results match the steady-state and asymptotic behavior in [35, 9]. The queue length concentration bound results are formally stated in Corollary 5.
@@
  • Probability of having an idle server: We obtain a bound on the probability of having an idle server at time $t$ (i.e. the queue length is less than $n$). Moreover, we also recover that as $n$ goes to infinity, this probability goes to 0 in the Super-Halfin-Whitt regime and goes to a value close to $1$ in the Sub-Halfin-Whitt regime. The full result is stated in Corollary 6.
@@
In summary, our finite-time behavior characterizations match the steady-state and asymptotic results previously established in [35, 9] and provide a holistic insight for Erlang-C systems.
1.2 Our approach: The Lyapunov-Poincaré method
In contrast to finite Markov chains, continuous state space Markov chains and Markov diffusions where there is a vast body of mixing literature for these settings [40, 53, 26, 65, 64, 3], there are only a few toolkits to analyze countable state space Markov chains due to a lack chain rule and many methods in the finite state space are no longer applicable due to the infiniteness of the state space. To this end, we aim to develop toolkits to analyze such systems using the Lyapunov drift, which exists when the system is positive recurrent.
Previously, it is well-known that one can obtain exponential ergodicity for a stochastic system if the system admits some Lyapunov drift outside of some finite set [51, 23, 68, 44, 5, 22, 50, 71, 2]. While one can easily get a handle on the mixing outside of the said set using the negative drift, it is much harder to handle the behavior inside the finite set. The special case when the finite set is a singleton (which is the case in an $M/M/1$ queue) is well understood [44]. When the finite set is not a singleton, one typically uses a ”stitching theorem” to combine knowledge of the drift outside the finite set and local mixing behavior inside the finite set to obtain a mixing time bound for the entire system. Prior work has explored the use of minorization [23, 68, 5, 71], contraction [26, 31] within the finite set or obtain the local Poincaré constant by using the radius of the finite set [21, 66, 20], and then a corresponding stitching theorem to obtain mixing rates. However, these approaches either do not lead to a tight mixing rate inside the finite set (especially for high-dimensional settings [63]) or are hard to obtain in practice. To this end, we propose the Lyapunov-Poincaré method, where we combine the Lyapunov drift outside of a finite set with a local Poincaré inequality inside the finite set to establish mixing results. Our work extends the previously established Lyapunov drift framework [51, 23, 68, 44, 5, 22, 50, 66, 71, 2] to allow a more fine-grained mixing analysis, which eventually enables us to obtain tight mixing bounds. In particular, we develop new stitching theorems so that depending on the properties of the finite set, we can better combine the drift properties outside the finite set with the local mixing behavior inside the finite set (which is characterized in terms of local Poincaré inequalities) so as to obtain more fine-grained mixing analysis.
  • Local canonical path method: The canonical path method, previously developed in [34], is a powerful technique used to analyze the mixing time of finite discrete Markov chains by bounding how ”congested” the state space is with respect to moving from one state to another. In particular, the canonical path method estimates the mixing time of a finite discrete Markov chain with the congestion ratio, which scales with the stationary distribution and the sum of the path length of paths passing through a particular edge. As such, the canonical path method can be viewed as a conductance-typed approach to obtain mixing. Notably, the canonical path method obtains a tight mixing bound when the Markov chain is a symmetric random walk on a path graph [6, 40], which is also the behavior of the Erlang-C system in the finite set when an appropriate Lyapunov function is chosen. By applying the local canonical path method, we obtain a tight weighted local Poincare inequality of the Erlang-C system inside the finite set, which allows us to use our stitching theorems to obtain the mixing rate bound for the entire system. We formally state our local canonical path lemma in Lemma 6.
  • Truncation: On the other hand, the aforementioned local canonical path method is only viable when the distribution inside the finite set is roughly uniform. For other finite sets whose distributions are not necessarily uniform, we propose the truncation method to obtain local mixing results inside such sets. In essence, if we can find another system with known Poincaré constant, which behaves identically inside the finite set, then we can use it to get a handle on the local Poincaré constant of the original system. With this method, we obtain tight mixing time bounds up to a constant factor for $M/M/n$ queues in the Mean Field regime, i.e. when $\lim_{n\rightarrow\infty}\frac{\lambda_{n}}{n}=c\in[0,1)$. For $c=0$ (the so-called Light Traffic regime), we have the mixing rate approaches to $1$, which matches the limiting behavior of the system.
In summary, our results and approach are summarized in Table 1 here.
Regime $M/M/\infty$ Mean Field Sub-HW Halfin-Whitt Super-HW $M/M/1$
@@ 1.2 Our approach: The Lyapunov-Poincaré method
$\lambda_{n}/n\rightarrow c$
V(q) Linear Constant + Exp Linear + Exp Two-sided Exp Exp
Finite set Singleton Truncation method Canonical path method Singleton
Mixing rate 1 $L_{n}$ $D_{n}$ $\Omega(1)$ $C_{n}(\sqrt{n}-\sqrt{\lambda_{n}})^{2}$ $(\sqrt{n}-\sqrt{\lambda_{n}})^{2}$
$e^{-\square t}$ $L_{n}\rightarrow 1-c$ $D_{n}\rightarrow\frac{1}{25}$ $C_{n}\rightarrow 1$
@@
Table 1: Overview of mixing time analysis using the Lyapunov-Poincaré method for $M/M/n$ for $\lambda_{n}=n-n^{1-\alpha},\mu=1$ where $\alpha$ is a parameter and $V(q)$ is the Lyapunov function. The number of waiting jobs is denoted as $\left[q-n\right]^{+}=\max\{q-n,0\}$, and the asymptotic behavior is the queue length distribution behavior as $n\rightarrow\infty$. When $\alpha\approx\infty$, this is called the Classical Heavy Traffic regime. For the Mean Field regime, we write $\alpha\approx 0$ as an intuitive explanation, whereas the actual condition is $\lim_{n\rightarrow\infty}\frac{\lambda_{n}}{n}=c\in[0,1)$.
1.3 Related works